OpenStudy (anonymous):

Gah, I accidentally logged out. The question I had posted was: An ambulance siren creates a 1200 Hz wave. What frequency will be heard if: The ambulance is moving away from you at 35 m/s? Toward you at 25 m/s?

OpenStudy (anonymous):

From the equations of Doppler effect, f'=(v/v+u)f if the source is moving away from you & f'=(v/v-u)f if it moves towards you v=speed of sound u=speed of the source f=actual frequency f'=perceived frequency

OpenStudy (anonymous):

so the answers are 1088 Hz & 1337.7 Hz respectively (JUST FOR CHECKING)